OpenStudy (ellamoyseyuk):

9. If a $2 bet is placed for a chance to win $500 for drawing a 10 of any suit from a standard deck of cards, what is the value of expectation?

OpenStudy (ellamoyseyuk):

@jim_thompson5910 can u help me

OpenStudy (agent0smith):

What's the probability of drawing a 10 from a deck of cards?

OpenStudy (agent0smith):

...how many 10's are in a deck of cards?

OpenStudy (ellamoyseyuk):

4

OpenStudy (agent0smith):

Now go back and answer the question

OpenStudy (ellamoyseyuk):

how would that answer my question

OpenStudy (agent0smith):

Because expectation is probability multiplied by the dollar outcome.

OpenStudy (agent0smith):

Why would i ask you something unrelated...

OpenStudy (ellamoyseyuk):

4 x 2 = 8

OpenStudy (agent0smith):

What's the probability of drawing a 10 from a deck of cards?

OpenStudy (agent0smith):

How many 10's are there in a deck. How many cards are there in a deck.

OpenStudy (ellamoyseyuk):

4/52 = 1/8

OpenStudy (agent0smith):

Incorrect simplifying... try again. What's the probability of NOT drawing a 10?

OpenStudy (ellamoyseyuk):

4/52 = 1/13

OpenStudy (ellamoyseyuk):

48/52 = 12/13

OpenStudy (agent0smith):

Expectation will be: probability of a 10 times profit, plus, probability of NOT a 10 times profit

OpenStudy (ellamoyseyuk):

is the answer 125

OpenStudy (agent0smith):

Idk. Show work. probability of a 10, times profit from drawing a 10 = ? probability of NOT a 10, times profit from NOT drawing a 10 = ?

OpenStudy (agent0smith):

If you bet $2 and a 10 is drawn, you get $500, for a profit of $498 If you don't draw a 10, you lose $2, for a profit of -$2